OpenStudy (anonymous):

A 2000 kg car experiences a constant braking force of 10,000 N, and comes to be a stop in 6 sec. What would be the distance traveled by the car in coming to rest?

OpenStudy (anonymous):

so f=ma

OpenStudy (anonymous):

and \[a=\frac{ \Delta v }{ \Delta t }\]

OpenStudy (anonymous):

How do I relate this to distance?

OpenStudy (anonymous):

so we know that our v final is 0

OpenStudy (anonymous):

and so we have have for our equation now\[10,000=2,000*\frac{ v-0 }{ 6 }\] you want to solve for v

OpenStudy (anonymous):

and then we can use the equation for one dimensional to solve for \[\Delta x\]

OpenStudy (anonymous):

that equation would be \[\Delta x=\frac{ 1 }{ 2 }(v+v _{0})t\]

OpenStudy (anonymous):

So the answer is 90 m?

OpenStudy (anonymous):

I got the same thing when I plugged it in, so I would say so

OpenStudy (anonymous):

or wait i made a slight error

OpenStudy (anonymous):

but it does not affect the out put

OpenStudy (anonymous):

Thank you for the help :)

OpenStudy (anonymous):

it is just the placement of the zero before so it should be\[10,000=2,000*\frac{ 0-v _{0} }{ 6 }\]

OpenStudy (anonymous):

you are welcome